OpenStudy (anonymous):

What is the simplified form of the expression 8sqrtv^2-sqrt^4-sqrt 9 x^2+sqrt 16^4? 15v2 + 11v 11v2 + 15v 5v2 + 3v 3v2 + 5v

OpenStudy (anonymous):

plz type it correctly

OpenStudy (anonymous):

\[8\sqrt{v^2}-\sqrt{v^4}-\sqrt{9v^2}+\sqrt{16^4}\]

OpenStudy (anonymous):